1. The average time of a pacemaker is a distribution 16 years It is requested a) Probability that a person to whom a pacemaker has been implanted must be implanted another one before 20 years b) If the pacemaker is working correctly 5 years in a patient, & What is the probability that you have to change it before 25 years? exponential with average
Answer #1

a) Probability that a person to whom a pacemaker has been implanted must be implanted another one before 20 years

=P(X<20)=1-e-x/ =1-e-20/16 =0.7135

b)P(X<25|X>5)=P(X<20)=1-e-x/ =1-e-20/16 =0.7135